shift summary

Sandrine Laplace (laplace@hep.uchicago.edu)
Wed, 26 May 1999 14:21:46 -0500

Shift Summary for owl shift, 05/26/97

On Shift: Sandrine Laplace
Hogan Nguyen

Hours of beam delivered at all intensities: 0
Hours of beam delivered at nominal intensity: 0
Hours of 99 data (special runs not included): 0
Hours of special runs (laser scans, muon runs...): 6

Accounting of lost time:

DYC BUSY of crate 59 stops trigger often. This crate has DYC+CTI+TDC

RSK found that NIM Discriminator that sends CLEAR to DYC in crate 59 is bad.
Replaced. Crate 59 now O.K. at 75 khz L1. Occasional stream 5
unknown crate errors at 150 khz L1, where L1/L2 = 6.

Added scaler to crate 59. Get flood of stream 5 unknown crate errors at
any rate.

Hogan worked on DC1, DC4.

With cosmic configuration and all beam triggers enabled, VME memory of
stream 6 occasionally fills up. This is due to B11, which we turned off.
B11 should be prescaled more.

DYC BUSY was full, so that it stopped trigering. A kfrend fixed it up.

========================================================================
Runs ended this shift:
Run 13138 - long laser scan
Run 13141 - cosmic run (still running at the end of the shift)

Tapes brought to FCC this shift: 0

********************************************************************
* LAPLACE Sandrine *
* --------------- *
* Enrico Fermi Institute KTeV 5514,S Blackstone Apt 318*
* 5640, S Ellis ave. at Chicago, IL 60637 *
* Chicago, IL 60637, USA Fermilab USA *
* http://www.geocities.com/Nashville/Rodeo/1190 *
********************************************************************